Sir, Here is my question. Q.A spring block system(2 blocks and a spring) has two blocks of mass 6kg and 4kg.Two forces of 20N and 10N acts on the 6kg and 4kg respectively in the opposite direction(away from the spring).When both the blocks have same accelaration what is the tension in the spring?

Sir,


Here is my question.


Q.A spring block system(2 blocks and a spring) has two blocks of mass 6kg and 4kg.Two forces of 20N and 10N acts on the 6kg and 4kg respectively in the opposite direction(away from the spring).When both the blocks have same accelaration what is the tension in the spring?

Grade:12

1 Answers

AskIITians Expert PRASAD IIT Kharagpur
14 years ago

Let T be the tension in the spring.

both blocks have same accelerarion i.e both blocks move in same direction with same magnitude of acceleration.

so,

a = (20 - T) / 6  = (T - 10) / 4

Solving above we get,

 

T = 14 N.